Q. A spherical tank of 1.2m radius is half filled with oil of relative density 0.8. Ifthe tank is given a horizontal acceleration of 10 m/s^2. Calculate the maximum pressure on the tank.
Grade:10

2 Answers

Aman Bansal
12 years ago

Dear Apurv,

The liquid rises along one of the sides. 

so,

Let the rise be h.

So,


h/2r=a/g

=> h= a/g * 2r

Since a=10m/s2

h=2r=2.4m

Therefore maximum pressure on the tank = ρgh
                                                         = 800 x 10 x 2.4
                                                         = 1.92 x 104 Pa

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Aditya
6 years ago
You assume that the base is passing through the centre to write h/2r...which beats the fact that the tank is half full